suppose that 16% of the population of the US is left-handed. If a random sample of 240 people from the US is chosen, approximate The probability that more than 39 or left-handed. Use the normal approximation to the by no meal with the correction for continually. Round your answer to that lease three decimal places do not round any intermediate step
Answer #1

Given data, Pa 16% => 0.16 F. P=0.16 q1-P q21-0.16 q= 0.84 1:0=0.84) Random Sample (n=240 (x)= 39 kle halle, P(X) = Nexopta p(240-39) 0.84 plan) = 206,92043] [9.4200-50] E 5.15337€ -43C . 1. p(39) : 5.15337€ -430